OpenStudy (anonymous):

Let f(x) = 8x3 – 28x + 61 and g(x) = 2x + 5. Find .

OpenStudy (anonymous):

is that suppose to be x to the third ?

OpenStudy (anonymous):

find f(g(x)) f(2x + 5) = 8(2x + 5)^3 -28(2x + 5) + 61

OpenStudy (anonymous):

= 8(2x + 5)(2x + 5)(2x + 5) - 28(2x + 5) +61